Chlorine is widely used to purify municipal water supplies and to treat swimming pool water. Suppose that the volume of a particular sample of Cl2 gas is 8.70 L at 890 torr and 24 ∘C.

1. How many grams of Cl2 are in the sample?

2. What volume will the Cl2 occupy at STP?

3. At what temperature will the volume be 15.50 L if the pressure is 875 torr?

4. At what pressure will the volume equal 7.00 L if the temperature is 55 ∘C?
Answer #1

1. PV = nRT . n = PV RT P = 890 tonn = 760 - 890 atm V = 8.70L T=(24+273)k = 297K R = 0.0821 lit atm motky in = 890 X 8.70 7630 PIVI = P V2 . Ti Te P = 890 tonn P = 875 toan. V = 8.7 lit V2 = 15.50 Lit T = 297K T = ? P₂ V₂ TL PVL - 875 X 15.50 X 297